### v4

# On the adjoint Selmer groups of semi-stable elliptic curves and Flach’s zeta elements

Abstract.
We explicitly construct the rank one primitive Stark (equivalently, Kolyvagin) system extending a constant multiple of Flach’s zeta elements for semi-stable elliptic curves. As its arithmetic applications, we obtain the equivalence between a specific behavior of the Stark system and the minimal modularity lifting theorem, and we also discuss the cyclicity of the adjoint Selmer groups. Our Stark system construction yields a more refined interpretation of the collection of Flach’s zeta elements than the “geometric Euler system” approach due to Flach, Wiles, Mazur, and Weston.
Key words and phrases:
Stark systems, symmetric squares of elliptic curves, Flach’s zeta elements, modularity lifting theorem, adjoint Selmer groups, refined Iwasawa theory
2020 Mathematics Subject Classification:
11R23, 11R39 (Primary); 11F33, 11F67,11G40 (Secondary)
[…]
Introduction
Flach’s zeta elements
As explained in Introduction of his celebrated work [Wil95], A. Wiles first tried to prove the semi-stable Shimura–Taniyama conjecture and Fermat’s last theorem via the construction of an Euler system for symmetric squares of semi-stable elliptic curves extending the cohomology classes constructed by M. Flach [Fla92], which we call Flach’s zeta elements in this article. However, the construction of the Euler system was incomplete, and it was regarded as a gap.
After that, Wiles, together with R. Taylor, found a completely different approach towards the semi-stable Shimura–Taniyama conjecture, which is now known as the Taylor–Wiles system argument [TW95], to fill the gap. See [Kur93, Kol94, RS94, Dar95] for the survey articles written before the gap was filled, and see [Cal23] for the far-reaching development of the Taylor–Wiles system argument.
This Euler system approach is often known as the “geometric Euler system” strategy, and it was further developed mainly by B. Mazur and T. Weston [Maz, MW, Wes01, Wes02]. However, it was still incomplete to obtain the exact bound of Selmer groups.
The main goals of this article are to investigate the behavior of the collection of Flach’s zeta elements beyond the scope of this geometric Euler system strategy and to deduce certain structural results for adjoint Selmer groups. However, it is well-known that the construction question of an Euler system extending Flach’s zeta element is extremely difficult (see [Rub00, §$\S$3.6], [Wes01, pp. 359–360], [MR04, Rem. 6.3.4], and [LZ19, Intro.] for example).
What is the new idea to overcome this notorious difficulty? We are greatly benefited from various remarkable and significant developments of the theory of Euler systems. In particular, it turns out that Kolyvagin systems are more essential than Euler systems to obtain the exact bound of the corresponding dual Selmer groups [MR04]. Also, the notion of Stark systems is developed by generalizing the units predicted by Stark-type conjectures [MR16]. The key observation is that the collection of Flach’s zeta elements forms a “Stark system of Gauss sum type” [^1] No Gauss sums actually appear in this article. Since we initially have only finitely many cohomology classes forming a Stark system, we call them “Gauss sum type”.. Since the module of Stark systems (equivalently, Kolyvagin systems) is free of rank one, the Gauss sum type Stark system naturally extends to the bona fide one. In other words, we construct derivative (so ramified) classes defined over $\mathbb{Q}$ directly from Flach’s zeta elements by using the framework of Stark systems.
Since there is a certain constant multiple subtlety in our Stark system approach, our result does not give a new proof of the semi-stable Shimura–Taniyama conjecture or Fermat’s last theorem yet. Also, we do not know whether our Stark system actually comes from an Euler system [^2] We are not sure whether it is possible to recover an Euler system from the Kolyvagin system associated to the Euler system.. Indeed, we prove that the minimal modularity lifting theorem is equivalent to showing that the constant multiple is a $p$-adic unit.
The cyclicity of adjoint Selmer groups
The cyclicity question has long history in Iwasawa theory. This is just a quick summary of the preface of Hida’s book [Hid22a] and we strongly recommend the reader to read it.
K. Iwasawa himself was interested in the structure of various classical Iwasawa modules including the cyclicity over the Iwasawa algebra. In [Iwa69], he obtained the cyclicity of certain classical Iwasawa modules over the Iwasawa algebra (as well as the main conjecture) under the Kummer–Vandiver conjecture. Even without assuming the Kummer–Vandiver conjecture, he proposed several interesting questions on the cyclicity around 1980’s [Iwa88, Iwa14].
H. Hida explored the non-abelian analogue of this nature, and [Hid86] was written with this purpose in mind. In particular, he studied the cyclicity question of the adjoint Selmer groups over the universal deformation ring in a series of his papers[Hid, Hid20, Hid22b, Hid22a]. In [Hid22a, Chap. 7], he proved many cyclicity results for the adjoint Selmer groups of Artin representations. As far as we understand, the elliptic curve case seems quite open [Hid22a, pp. 287–288 and 339] unless the length of adjoint Selmer groups is $\leq 1$. As an application of the above Stark system construction, we obtain a cyclicity result for the adjoint Selmer groups of elliptic curves (Corollary 1.5).
Comparison with other recent work
Recent years, some groups of mathematicians have developed new approaches towards the construction of symmetric square and adjoint Euler systems from different sources [LZ19, Urb21, LZ, SS]. We provide brief remarks on their constructions and highlight the novelty of our approach in contrast to theirs, as per the request to address this aspect.
In [LZ19], an Euler system is constructed from Beilinson–Flach’s elements and bounds Selmer groups of the symmetric square representations of an ordinary modular form twisted by a non-trivial Dirichlet character. This twist is inevitable to have the correct Euler system relation. In [Urb21], an Euler system for adjoint representations of ordinary Hilbert modular forms is constructed directly from congruence modules, but the $R=\mathbb{T}$ theorem is an ingredient of the construction. Also, the corresponding explicit reciprocity law depends on a conjecture on the Fitting ideals of some equivariant congruence modules for abelian base change. In [LZ], an Euler system for adjoint representations of ordinary modular forms is constructed from Asai–Flach classes and is used to prove one-sided divisibility of the main conjecture for the symmetric square Selmer group over the cyclotomic $\mathbb{Z}_{p}$-extension of $\mathbb{Q}$ for all but finitely many height one prime ideals.
In [SS], an Euler system for the symmetric square representation of an ordinary modular form twisted by an odd Dirichlet character is initially constructed by utilizing the pull-back of Eisenstein series on $\mathrm{SO}(3,2)$. Then by using Hida theory (among others), an Euler system for the symmetric square representation (with no extra twist) is obtained. As its consequence, the $R=\mathbb{T}$ theorem follows. The $p$-ordinary condition is assumed in all the work mentioned above.
Our Stark system also detects the precise upper bound of the adjoint Selmer groups, so the direct connection with the modularity lifting theorem is obtained. More precisely, our construction yields an equivalence statement as mentioned earlier. It appears that our approach is the most akin to Wiles’ original method since we work with Flach’s zeta elements as Wiles did, so it might have some historical meaning. Also, our argument works for any good reduction prime. In addition, the structural refinement like the cyclicity is not observed in other approaches.
The organization
In §1, we state our main result on the existence of the rank one primitive Stark systems for symmetric square representations and its arithmetic applications. In §2, we review Flach’s zeta elements and state the explicit reciprocity law for Flach’s zeta elements. In §3, we verify the working hypotheses for rank one Stark systems. In §4, we recall Weston’s computation on the explicit reciprocity law. In §5, we prove our main result by constructing the Stark system explicitly. This is the heart of this article. In §6, we prove corollaries to the main result.
1. Statement of the main result
Let $E$ be a semi-stable modular [^3] We do not use the semi-stable modularity theorem of Wiles and Taylor–Wiles [Wil95, TW95] and its generalizations in this article. elliptic curve over $\mathbb{Q}$ of conductor $N$ and $p\geq 5$ a good reduction prime for $E$. Denote by $T_{p}E$ the $p$-adic Tate module for $E$. Assume that $\overline{\rho}:G_{\mathbb{Q}}=\mathrm{Gal}(\overline{\mathbb{Q}}/\mathbb{Q})\to\mathrm{Aut}_{\mathbb{F}_{p}}(E[p])$ is irreducible and $N=N(\overline{\rho})$ where $N(\overline{\rho})$ is the conductor of $\overline{\rho}$.
The main goal of this article is to construct the (rank one) Stark system from Flach’s zeta elements. It partially resolves a question of Mazur–Rubin in [MR04, Rem. 6.3.4]. Indeed, we completely settle their question under the semi-stable modularity lifting theorem (see Corollary 1.4 below).
Theorem 1.1 (Flach–Stark systems).
For a prime $\ell$ with $(Np,\ell)=1$, $\ell\equiv 1\pmod{p}$, and $a_{\ell}(E)\not\equiv\pm 2\pmod{p}$, there exists the non-trivial primitive rank one Stark system
$\boldsymbol{\epsilon}^{\mathrm{Flach},\ell}$
for $\mathrm{Sym}^{2}(T_{p}E)$ extending a constant multiple of Flach’s zeta element $c(\ell)$ (recalled in Theorem 2.2) which determines the structure of the $\ell$-strict (trace zero) adjoint Bloch–Kato Selmer group
$\mathrm{Sel}_{\ell\textrm{-}\mathrm{str}}(\mathbb{Q},\mathrm{ad}^{0}(E[p^{\infty}])).$
In particular, the constant multiple is independent of $\ell$. We call $\boldsymbol{\epsilon}^{\mathrm{Flach},\ell}$ the Flach–Stark system.
Proof.
The explicit construction of $\boldsymbol{\epsilon}^{\mathrm{Flach},\ell}$ is given in §5. ∎
In the rest of this section, we fix a prime $\ell$ satisfying the conditions in Theorem 1.1. Let $\boldsymbol{\kappa}^{\mathrm{Flach},\ell}$ be the (non-trivial and primitive) Flach–Kolyvagin system corresponding to $\boldsymbol{\epsilon}^{\mathrm{Flach},\ell}$ under the isomorphism between the modules of Stark systems and Kolyvagin systems under our setting [MR16, Rem. 11.5 and Thm. 12.4]. Indeed, the constant multiple in Theorem 1.1 is designed to deduce the following formula for the exact bound of the genuine Bloch–Kato Selmer group in terms of $\boldsymbol{\kappa}^{\mathrm{Flach},\ell}$.
Corollary 1.2 (Exact bound).
Let
$\mathrm{loc}^{s}_{\ell}(\boldsymbol{\kappa}^{\mathrm{Flach},\ell})=\left\{\mathrm{loc}^{s}_{\ell}(\kappa^{\mathrm{Flach},\ell}_{n}):n\in\mathcal{N}^{\mathrm{Flach},(\ell)}_{1}\right\}$
be the singular quotient of the localization of $\boldsymbol{\kappa}^{\mathrm{Flach},\ell}$ at $\ell$ where $\mathcal{N}^{\mathrm{Flach},(\ell)}_{1}$ is defined in Definition 2.3. Then
$\mathrm{ord}_{p}(\mathrm{loc}^{s}_{\ell}(\kappa^{\mathrm{Flach},\ell}_{1}))=\mathrm{length}_{\mathbb{Z}_{p}}(\mathrm{Sel}(\mathbb{Q},\mathrm{ad}^{0}(E[p^{\infty}])))$
where $\mathrm{Sel}(\mathbb{Q},\mathrm{ad}^{0}(E[p^{\infty}]))$ is the adjoint Bloch–Kato Selmer group and $\mathrm{ord}_{p}(\mathrm{loc}^{s}_{\ell}(\kappa^{\mathrm{Flach},\ell}_{1}))$ is the $p$-divisibility index of $\mathrm{loc}^{s}_{\ell}(\kappa^{\mathrm{Flach},\ell}_{1})$ in $\mathrm{H}^{1}_{/f}(\mathbb{Q}_{\ell},\mathrm{Sym}^{2}(T_{p}E))$.
Proof.
See §$\S$6.1. ∎
Remark 1.3.
It is known that the primitivity of a (rank one) Kolyvagin system is closely related to showing the corresponding main conjecture [MR04, KKS20, Sak22, Kim]. This principle would apply to our setting as follows if such a formulation is valid. Let $\mathbb{Q}_{\infty}$ be the cyclotomic $\mathbb{Z}_{p}$-extension of $\mathbb{Q}$ and $\Lambda$ the cyclotomic Iwasawa algebra. If the $\Lambda$-adic deformation $\boldsymbol{\kappa}^{\mathrm{Flach},\ell,\infty}$ of $\boldsymbol{\kappa}^{\mathrm{Flach},\ell}$ exists and $\kappa^{\mathrm{Flach},\ell,\infty}_{1}$ is non-zero in $\widehat{\mathrm{Sel}_{\ell\textrm{-}\mathrm{rel}}}(\mathbb{Q}_{\infty},\mathrm{Sym}^{2}(T_{p}E))$, then the hypothetical main conjecture
``$\mathrm{char}_{\Lambda}\left(\widehat{\mathrm{Sel}_{\ell\textrm{-}\mathrm{rel}}}(\mathbb{Q}_{\infty},\mathrm{Sym}^{2}(T_{p}E))/\Lambda\kappa^{\mathrm{Flach},\ell,\infty}_{1}\right)=\mathrm{char}_{\Lambda}\left(\mathrm{Sel}_{\ell\textrm{-}\mathrm{str}}(\mathbb{Q}_{\infty},\mathrm{ad}^{0}(E[p^{\infty}]))^{\vee}\right)"$
[…]
Denote by $\mathrm{deg}(\phi)$ the minimal modular degree of $E$. We obtain the following connection between Flach’s zeta elements and the minimal modularity lifting theorem from Theorem 1.1.
Corollary 1.4 (Minimal modularity lifting).
The following three statements are equivalent:
  1. (1)
    The constant multiple in Theorem 1.1 is a $p$-adic unit so that
    $\mathrm{ord}_{p}(\mathrm{loc}^{s}_{\ell}c(\ell))=\mathrm{length}_{\mathbb{Z}_{p}}\mathrm{Sel}(\mathbb{Q},\mathrm{ad}^{0}(E[p^{\infty}])).$
  2. (2)
    The Bloch–Kato conjecture for the adjoint representation of $E$ holds, i.e.
    $\mathrm{ord}_{p}(\mathrm{deg}(\phi))=\mathrm{length}_{\mathbb{Z}_{p}}\mathrm{Sel}(\mathbb{Q},\mathrm{ad}^{0}(E[p^{\infty}])).$
  3. (3)
    The minimal deformation ring of $\overline{\rho}$ and the minimal Hecke algebra localized at the maximal ideal $\mathfrak{m}_{\overline{\rho}}$ associated to $\overline{\rho}$ are isomorphic as complete intersections, i.e.
    $R^{\mathrm{min}}_{\overline{\rho}}\simeq\mathbb{T}^{\mathrm{min}}_{\mathfrak{m}_{\overline{\rho}}}.$
Proof.
See §$\S$6.2. ∎
Of course, all the statements in Corollary 1.4 are true thanks to the work of Wiles and Taylor–Wiles [Wil95, TW95]. One remarkable feature of Corollary 1.4 is that the modularity lifting theorem can be deduced from a certain property regarding Flach’s zeta elements (finally!).
On the other hand, the modularity lifting theorem (Corollary 1.4.(3)) yields the explicit construction of the rank one primitive Kolyvagin system for $\mathrm{Sym}^{2}(T_{p}E)$ extending Flach’s zeta elements (without introducing any constant multiple) for every elliptic curve $E$ which occurs in the minimal deformation space of $\overline{\rho}$ [^4] We thank H. Hida for bringing this deformation-theoretic observation to our attention. .
From the viewpoint of Corollary 1.4, we anticipate that our construction of the Stark system will contribute directly to the revival of Wiles’ abandoned “geometric Euler system” approach to the semi-stable Shimura–Taniyama conjecture in a future.
Although we do not know whether or how $\mathrm{loc}^{s}_{\ell}(\boldsymbol{\kappa}^{\mathrm{Flach},\ell})$ determines the structure of $\mathrm{Sel}(\mathbb{Q},\mathrm{ad}^{0}(E[p^{\infty}]))$ in general yet [^5] In [Kim], the self-duality of Galois representations is used in an essential way. We are informed that R. Sakamoto is working on the general structure theorem with rank zero Stark/Kolyvagin systems, and it seems to be closely related to our approach., we still have the following structural application under the $p$-indivisibility assumption on $c(\ell)$.
Corollary 1.5 (Cyclicity).
Suppose that any (equivalent) statement of Corollary 1.4 holds. If we further assume that $c(\ell)\in\mathrm{Sel}_{\ell\textrm{-}\mathrm{rel}}(\mathbb{Q},\mathrm{Sym}^{2}(T_{p}E))$ is not divisible by $p$, then the adjoint Selmer group is cyclic, i.e.
$\mathrm{Sel}(\mathbb{Q},\mathrm{ad}^{0}(E[p^{\infty}]))\simeq\mathbb{Z}_{p}/\mathrm{deg}(\phi)\mathbb{Z}_{p}$
where $\mathrm{Sel}_{\ell\textrm{-}\mathrm{rel}}(\mathbb{Q},\mathrm{Sym}^{2}(T_{p}E))$ is the $\ell$-relaxed compact symmetric square Bloch–Kato Selmer group.
Proof.
See §$\S$6.3. ∎
This cyclicity can be understood as the structural refinement of the Bloch–Kato conjecture (cf. [Kim, Kim24]). We make the following question (not conjecture).
Question 1.6.
Does the $p$-indivisibility assumption in Corollary 1.5 hold in general? It is equivalent to proving that $\mathrm{Sel}_{\ell\textrm{-}\mathrm{str}}(\mathbb{Q},\mathrm{ad}^{0}(E[p^{\infty}]))$ is trivial.
The higher weight generalization of Theorem 1.1 and its interpretation in terms of rank zero Stark/Kolyvagin systems will be investigated in a joint work with Ryotaro Sakamoto.
2. Flach’s zeta elements and the explicit reciprocity law
Let $G_{k}\subseteq\mathrm{GL}_{2}(\mathbb{Z}/p^{k}\mathbb{Z})$ be the image of $G_{\mathbb{Q}}$ under the mod $p^{k}$ reduction of $\rho:G_{\mathbb{Q}}\to\mathrm{Aut}_{\mathbb{Z}_{p}}(T_{p}E)\simeq\mathrm{GL}_{2}(\mathbb{Z}_{p})$. The following annihilation result is the very starting point of this story.
Theorem 2.1 (Flach).
Assume that $p$ is odd, $\overline{\rho}$ is absolutely irreducible, $p$ does not divide $N$, and $N=N(\overline{\rho})$. If we further assume that $N$ is square-free and $\mathrm{H}^{1}(G_{k},\mathrm{ad}^{0}(E[p^{k}]))=0$ for all $k\geq 1$, then
$\mathrm{deg}(\phi)\cdot\mathrm{Sel}(\mathbb{Q},\mathrm{ad}^{0}(E[p^{\infty}]))=0.$
In particular, $\mathrm{Sel}(\mathbb{Q},\mathrm{ad}^{0}(E[p^{\infty}]))$ is finite, and it is trivial for all but finitely many primes $p$.
Proof.
See [Fla95, Thm. 1] and [Fla92, Thm. 1]. ∎
In order to obtain Theorem 2.1, Flach constructed the following important cohomology classes, which we call Flach’s zeta elements.
Theorem 2.2 (Flach’s zeta elements).
Assume that $p$ is odd, $\overline{\rho}$ is absolutely irreducible, $p$ does not divide $N$, and $N=N(\overline{\rho})$. For a prime $\ell$ not dividing $2Np$, there exists a cohomology class
$c(\ell)\in\mathrm{Sel}_{\ell\textrm{-}\mathrm{rel}}(\mathbb{Q},\mathrm{Sym}^{2}(T_{p}E))$
such that
  • $\mathrm{loc}_{q}(c(\ell))\in\mathrm{H}^{1}_{f}(\mathbb{Q}_{q},\mathrm{Sym}^{2}(T_{p}E))$ for any prime $q\neq\ell$, and
  • $\mathrm{ord}_{p}(\mathrm{loc}^{s}_{\ell}(c(\ell)))\leq\mathrm{ord}_{p}(\mathrm{deg}(\phi))+\mathrm{ord}_{p}(\alpha_{\ell}-\beta_{\ell})$
where $\alpha_{\ell}$ and $\beta_{\ell}$ are the roots of $X^{2}-a_{\ell}(E)X+\ell$.
Proof.
See [Fla95, Prop. 1 and §$\S$5]. ∎
Definition 2.3.
Let $k\geq 1$ be an integer. A prime $\ell$ is said to be a $k$-Flach prime if $(\ell,Np)=1$, $\ell\equiv 1\pmod{p^{k}}$, and $a_{\ell}(E)\not\equiv\pm 2\pmod{p}$. This condition is equivalent to that $(\ell,Np)=1$ and the arithmetic Frobenius at $\ell$ acts on $E[p^{k}]$ as $\tau$ chosen in Choice 3.4 below. We call 1-Flach primes by Flach primes for convenience. Denote by $\mathcal{P}^{\mathrm{Flach},(\ell)}_{k}$ the set of all $k$-Flach primes except $\ell$ for fixed $(E,p)$ and by $\mathcal{N}^{\mathrm{Flach},(\ell)}_{k}$ the set of square-free products of primes in $\mathcal{P}^{\mathrm{Flach},(\ell)}_{k}$.
It is easy to check that $\alpha_{\ell}-\beta_{\ell}$ is a $p$-adic unit for a Flach prime $\ell$. From Theorem 2.2, we have inequality
$\mathrm{ord}_{p}(\mathrm{loc}^{s}_{\ell}(c(\ell)))\leq\mathrm{ord}_{p}(\mathrm{deg}(\phi))$
for a Flach prime $\ell$. Here, $\mathrm{deg}(\phi)$ plays the role of the depth of a partial geometric Euler system in the sense of Mazur and Weston [Maz, Wes01, Wes02].
This inequality can be upgraded to equality thanks to Weston’s computation [Wes01, Wes02]. The equality (2.1) below should be viewed as the explicit reciprocity law for Flach’s zeta elements since $\mathrm{deg}(\phi)$ has a precise connection with the adjoint $L$-value via the formula of Shimura and Hida.
Proposition 2.4 (Weston’s explicit reciprocity law).
If $\ell$ is a Flach prime, the inequality in Theorem 2.2 becomes equality
(2.1) $\mathrm{ord}_{p}(\mathrm{loc}^{s}_{\ell}c(\ell))=\mathrm{ord}_{p}(\mathrm{deg}(\phi)).$
Proof.
See §$\S$4. ∎
The collection of $c(\ell)$ varying Flach primes $\ell$ forms a (cohesive) Flach system of depth $\mathrm{deg}(\phi)$ in the sense of Mazur and Weston [Maz, Wes02].
3. Verifying the hypotheses for Stark systems
We review the running hypotheses of Stark systems (of core rank one) in [MR16, §$\S$4] and verify them for our setting.
3.1. Running hypotheses
Let $\Sigma=\{p,\infty,\textrm{ramified primes for }T_{p}E\}$ and fix a Flach prime $\ell$.
Definition 3.1.
By Selmer data, we mean the triple
$(\mathrm{Sym}^{2}(T_{p}E),\mathcal{F}^{\ell}_{\mathrm{BK}},\mathcal{P}^{\mathrm{Flach},(\ell)})$
where
  • $\mathrm{Sym}^{2}(T_{p}E)$ as a $G_{\mathbb{Q}}$-module,
  • $\mathcal{F}^{\ell}_{\mathrm{BK}}$ is the Bloch–Kato Selmer structure except the relaxed local condition at $\ell$, and
  • $\mathcal{P}^{\mathrm{Flach},(\ell)}$ is the set of Flach primes excluding $\ell$.
For a $G_{\mathbb{Q}}$-module $A$, we write $\mathbb{Q}(A)$ for the fixed field in $\overline{\mathbb{Q}}$ of the kernel of the map $G_{\mathbb{Q}}\to\mathrm{Aut}(A)$.
Assumption 3.2.
We list the assumptions to proceed the Stark system argument [MR16, §$\S$4]:
  1. (H.1)
    $\mathrm{H}^{0}(\mathbb{Q},\mathrm{Sym}^{2}(E[p]))=\mathrm{H}^{0}(\mathbb{Q},\mathrm{ad}^{0}(E[p]))=0$ and $\mathrm{Sym}^{2}(E[p])$ is absolutely irreducible.
  2. (H.2)
    There exists an element $\tau\in G_{\mathbb{Q}(\zeta_{p^{\infty}})}=\mathrm{Gal}(\overline{\mathbb{Q}}/\mathbb{Q}(\zeta_{p^{\infty}}))$ such that $\mathrm{Sym}^{2}(T_{p}E)/(\tau-1)\mathrm{Sym}^{2}(T_{p}E)$ is free of rank one over $\mathbb{Z}_{p}$.
  3. (H.3)
    $\mathrm{H}^{1}(\mathbb{Q}_{T}/\mathbb{Q},\mathrm{Sym}^{2}(E[p]))=\mathrm{H}^{1}(\mathbb{Q}_{T}/\mathbb{Q},\mathrm{ad}^{0}(E[p]))=0$ where $\mathbb{Q}_{T}=\mathbb{Q}(\mathrm{Sym}^{2}(T_{p}E),\zeta_{p^{\infty}})$.
  4. (H.4)
    $\mathrm{Sym}^{2}(E[p])\not\simeq\mathrm{ad}^{0}(E[p])$ as Galois modules or $p>3$.
  5. (H.5)
    The Selmer structure $\mathcal{F}^{\ell}_{\mathrm{BK}}$ is cartesian.
  6. (H.6)
    $\chi(\mathrm{Sym}^{2}(T_{p}E),\mathcal{F}^{\ell}_{\mathrm{BK}})=1$, where $\chi(\mathrm{Sym}^{2}(T_{p}E),\mathcal{F}^{\ell}_{\mathrm{BK}})$ is the core rank of $(\mathrm{Sym}^{2}(T_{p}E),\mathcal{F}^{\ell}_{\mathrm{BK}})$.
  7. (H.7)
    $I_{q}=0$ for $q\in\mathcal{P}^{\mathrm{Flach},(\ell)}_{k}$ when the coefficient ring is $\mathbb{Z}/p^{k}\mathbb{Z}$.
The following lemma is useful.
Lemma 3.3.
Let $E$ be a semi-stable elliptic curve over $\mathbb{Q}$, $p$ a prime, and $\overline{\rho}:G_{\mathbb{Q}}\to\mathrm{Aut}_{\mathbb{F}_{p}}(E[p])$. Then $\overline{\rho}$ is surjective if and only if $\overline{\rho}$ is irreducible.
Proof.
See [Edi97, Prop. 2.1]. ∎
3.2. Verifying (H.1)
(H.1) follows from Lemma 3.3 and [Rub00, §$\S$3.6].
3.3. Verifying (H.2)
Thanks to Lemma 3.3, we are able to make the following choice of a Galois element.
Choice 3.4.
We choose an element $\tau\in G_{\mathbb{Q}(\zeta_{p^{\infty}})}$ such that
$\rho(\tau)\sim\begin{pmatrix}\alpha&0\\ 0&\alpha^{-1}\end{pmatrix}$
where $\alpha\in\mathbb{Z}_{p}$ with $\alpha^{2}\not\equiv 1\pmod{p}$ when $p>3$, or $\alpha\in\mathbb{Z}_{p^{2}}$ with $\alpha\equiv\sqrt{-1}\in\mathbb{F}_{3^{2}}$ when $p=3$.
This choice of $\tau$ is possible and satisfies (H.2) for $\mathrm{Sym}^{2}(T_{p}E)$ as explained in [Rub00, §$\S$3.6]. The choice of $\tau$ determines the type of auxiliary primes when we apply the Chebotarev density theorem as in Definition 2.3. In the case of Heegner points, $\tau$ is chosen to be the complex conjugation. In [Fla92], $\tau$ is chosen to be the complex conjugation again. For Kato’s Euler systems, $\tau$ is chosen to be a unipotent element as explained in [Rub00, Prop. 3.5.8]. We also recommend the reader to read [Wil95, Intro.] for his change of auxiliary primes.
3.4. Verifying (H.3)
(H.3) follows from [Rub00, §$\S$3.6]. See also [Fla92, Lem. 1.2] with [Fla95, Rem. 1 in §$\S$4], and [DDT97, Lem. 2.48] when $p>5$.
3.5. Verifying (H.4)
This is immediate.
3.6. Verifying (H.5)
(H.5) follows from [MR04, Lem. 3.7.1 and Rem. 3.7.2].
3.7. Verifying (H.6)
As explained in [MR04, Rem. 6.3.4], the core rank of Kolyvagin systems for $\left(\mathrm{Sym}^{2}(T_{p}E),\mathcal{F}^{\ell}_{\mathrm{BK}},\mathcal{P}^{\mathrm{Flach},(\ell)}\right)$ is one since $\ell$ is a Flach prime. The core rank of Stark systems is defined in the exactly same way. See also [MR04, Thm. 4.1.13] and [MR16, Prop. 3.3].
3.8. Verifying (H.7)
Suppose that we are working over the $\mathbb{Z}/p^{k}\mathbb{Z}$-coefficients. For $q\in\mathcal{P}^{\mathrm{Flach},(\ell)}_{k}$, $I_{q}$ is defined by the ideal of $\mathbb{Z}/p^{k}\mathbb{Z}$ generated by $(1-\alpha^{2}_{q})\cdot(1-q)\cdot(1-\beta^{2}_{q})$ and $(1-q)$. This means that $I_{q}=(1-q)=(0)$, so we are done.
4. A proof of the explicit reciprocity law
We recall the computation done by Weston in [Wes01] but with a very slight modification. Proposition 2.4 follows from this computation.
4.1. A lemma
Lemma 4.1.
Suppose that $q$ is prime to $Np$. Then $\mathrm{H}^{1}_{/f}(\mathbb{Q}_{q},\mathrm{Sym}^{2}(T_{p}E))\simeq\mathrm{H}^{0}(\mathbb{F}_{q},\mathrm{ad}^{0}(T_{p}E))$. If $q$ is a $k$-Flach prime, then we have isomorphism
$\psi_{q}:\mathrm{H}^{1}_{/f}(\mathbb{Q}_{q},\mathrm{Sym}^{2}(E[p^{k}]))\simeq\mathbb{Z}/p^{k}\mathbb{Z}.$
Proof.
See [Wes01, Lem. 10.9] for the first statement and the mod $p^{k}$ version also holds by the same reasoning. Since [Wes01, Lem. 10.9] concerns the complex conjugation, we give the detail for the latter. Since $q$ is a $k$-Flach prime, the arithmetic Frobenius $\mathrm{Fr}_{q}$ at $q$ acts on $E[p^{k}]$ by $\tau$ in Choice 3.4. Choose a basis $x$, $y$ of $E[p^{k}]$ such that $\mathrm{Fr}_{q}(x)=\alpha\cdot x$ and $\mathrm{Fr}_{q}(y)=\alpha^{-1}\cdot y$. Then $x\otimes x$, $x\otimes y+y\otimes x$, $y\otimes y$ forms a basis of $\mathrm{Sym}^{2}(E[p^{k}])$. Since $\mathrm{Sym}^{2}(E[p^{k}])(-1)=\mathrm{ad}^{0}(E[p^{k}])$, $\mathrm{Fr}_{q}$ acts on the induced basis of $\mathrm{ad}^{0}(E[p^{k}])$ by multiplication by $\alpha^{2}q^{-1}$, $q^{-1}$, and $\alpha^{-2}q^{-1}$, respectively. Since $q\equiv 1\pmod{p^{k}}$ and $\alpha^{2}\not\equiv 1\pmod{p}$, $\mathrm{H}^{1}_{/f}(\mathbb{Q}_{q},\mathrm{Sym}^{2}(E[p^{k}]))\simeq\mathrm{H}^{0}(\mathbb{F}_{q},\mathrm{ad}^{0}(E[p^{k}]))$ is free of rank one over $\mathbb{Z}/p^{k}\mathbb{Z}$, so we are done. ∎
4.2. The computation of the image
We recap Weston’s computation of the image of $c(\ell)$ under $\mathrm{loc}^{s}_{\ell}$ but with Flach prime $\ell$. This computation slightly refines [Fla95, Lem. 2.5], and see [Wes02, Thm. 3.1.1] for a more theoretical background of this computation.
As in Lemma 4.1, we choose a basis $x$, $y$ of $T_{p}E$ with respect to which $\mathrm{Fr}_{\ell}$ has matrix
$\begin{pmatrix}\alpha&0\\ 0&\alpha^{-1}\cdot\ell\end{pmatrix}$
where $\alpha\in\mathbb{Z}_{p}$ with $\alpha^{2}\not\equiv 1\pmod{p}$. By using the same idea of Lemma 4.1 again, we have isomorphism
$\mathrm{H}^{1}_{/f}(\mathbb{Q}_{\ell},\mathrm{Sym}^{2}(T_{p}E))$ $\simeq\mathrm{H}^{0}(\mathbb{F}_{\ell},\mathrm{ad}^{0}(T_{p}E))$
$=\mathbb{Z}_{p}\cdot\begin{pmatrix}1&0\\ 0&-1\end{pmatrix}.$
Following Weston’s computation in [Wes01, p. 369], the image of $\mathrm{loc}^{s}_{\ell}c(\ell)$ in $\mathrm{H}^{0}(\mathbb{F}_{\ell},\mathrm{ad}^{0}(T_{p}E))$ is
$6\cdot\mathrm{deg}(\phi)\cdot(\alpha_{\ell}-\beta_{\ell})\cdot\begin{pmatrix}1&0\\ 0&-1\end{pmatrix}.$
Since $p\geq 5$ and $\ell$ is a Flach prime, we have an isomorphism of cyclic modules
$\dfrac{\mathrm{H}^{1}_{/f}(\mathbb{Q}_{\ell},\mathrm{Sym}^{2}(T_{p}E))}{\mathbb{Z}_{p}\cdot\mathrm{loc}^{s}_{\ell}c(\ell)}\simeq\mathbb{Z}_{p}/\mathrm{deg}(\phi)\mathbb{Z}_{p},$
which implies the explicit reciprocity law (2.1).
5. The construction of Flach–Stark systems
We freely use the language in [MR04, MR16] in order to keep the argument concise.
5.1. Basic setup for Stark systems
Due to Theorem 2.1 and [MR04, Lem. 3.5.3], we are able to and do fix an integer $k\gg 0$ such that
$\mathrm{Sel}_{\mathcal{F}_{\mathrm{BK}}}(\mathbb{Q},\mathrm{ad}^{0}(E[p^{k}]))=\mathrm{Sel}_{\mathcal{F}_{\mathrm{BK}}}(\mathbb{Q},\mathrm{ad}^{0}(E[p^{\infty}]))$
where $\mathcal{F}_{\mathrm{BK}}$ is the Bloch–Kato Selmer structure.
Fix a Flach prime $\ell$ and let
(5.1) $n=q_{1}\cdot q_{2}\cdot\cdots\cdot q_{s}\in\mathcal{N}^{\mathrm{Flach},(\ell)}_{k}$
throughout this section.
Remark 5.1.
In the case of Kolyvagin systems of Gauss sums [MR04, Rem. 6.3.3], the choice of $n$ depends seriously on $\ell$. Namely, every prime divisor of $n$ should divide $\ell-1$. In our case, there are no such restrictions, so each $q_{i}$ is just a $k$-Flach prime not equal to $\ell$.
Following [MR16, §$\S$6], recall that
$W_{n}$ $=\bigoplus^{s}_{i=1}\mathrm{Hom}\left(\mathrm{H}^{1}_{/f}(\mathbb{Q}_{q_{i}},\mathrm{Sym}^{2}(E[p^{k}])),\mathbb{Z}/p^{k}\mathbb{Z}\right),$
$Y_{n}$ $=\bigwedge^{1+\nu(n)}\mathrm{Sel}_{\mathcal{F}^{\ell n}_{\mathrm{BK}}}(\mathbb{Q},\mathrm{Sym}^{2}(E[p^{k}]))\otimes\bigwedge^{\nu(n)}W_{n}$
where $\nu(n)=s$ is the number of prime divisors of $n$, and $\mathcal{F}^{\ell n}_{\mathrm{BK}}$ is the $\ell n$-relaxed Bloch–Kato Selmer structure. For $m$ dividing $n$, we have the cartesian square
[…]
and the canonical map $\Psi_{n,m}:Y_{n}\to Y_{m}$ attached to the above square as in [MR16, Prop. A.2]. For a $k$-Flach prime $q$, we have an isomorphism
$\mathrm{H}^{1}_{/f}(\mathbb{Q}_{q},\mathrm{Sym}^{2}(E[p^{k}]))\simeq\mathrm{H}^{1}_{\mathrm{tr}}(\mathbb{Q}_{q},\mathrm{Sym}^{2}(E[p^{k}]))$
by [MR04, Lem. 1.2.4] where $\mathrm{H}^{1}_{\mathrm{tr}}$ means the transverse local condition.
5.2. Definition and properties of Stark systems
The canonical map $\Psi_{n,m}$ given above has the following compatibility.
Proposition 5.2.
Suppose $n_{0}\in\mathcal{N}^{\mathrm{Flach},(\ell)}_{k}$, $n_{1}$ divides $n_{0}$, and $n_{2}$ divides $n_{1}$. Then
$\Psi_{n_{0},n_{2}}=\Psi_{n_{1},n_{2}}\circ\Psi_{n_{0},n_{1}}.$
Proof.
See [MR16, Prop. 6.4]. ∎
We briefly recall the definition of Stark systems [MR16, Def. 6.5] and their important properties.
Definition 5.3.
The $\mathbb{Z}/p^{k}\mathbb{Z}$-module of the (rank one) Stark system $\mathbf{SS}_{1}(\mathrm{Sym}^{2}(E[p^{k}]))=\mathbf{SS}_{1}(\mathrm{Sym}^{2}(E[p^{k}]),\mathcal{F}^{\ell}_{\mathrm{BK}},\mathcal{P}^{\mathrm{Flach},(\ell)}_{k})$ for Selmer data $(\mathrm{Sym}^{2}(E[p^{k}]),\mathcal{F}^{\ell}_{\mathrm{BK}},\mathcal{P}^{\mathrm{Flach},(\ell)}_{k})$ is defined to be the inverse limit
[…]
Theorem 5.4.
Under our working hypotheses, we have $\mathbf{SS}_{1}(\mathrm{Sym}^{2}(E[p^{k}]))\simeq\mathbb{Z}/p^{k}\mathbb{Z}$.
Proof.
See [MR16, Thm. 6.7]. ∎
A Stark system $\boldsymbol{\epsilon}^{\ell}$ is primitive if the image of $\boldsymbol{\epsilon}^{\ell}$ in $\mathbf{SS}_{1}(\mathrm{Sym}^{2}(E[p]))$ is non-zero [MR16, Prop. 7.3 and Thm. 7.4]. The theory of Stark systems yields the following result on the exact size of the dual Selmer groups.
Theorem 5.5.
We keep our working hypotheses, and let $\boldsymbol{\epsilon}^{\ell}$ be a Stark system.
  1. (1)
    If $\boldsymbol{\epsilon}^{\ell}$ is non-trivial, then
    $\mathrm{length}_{\mathbb{Z}_{p}}\mathrm{Sel}_{\ell\textrm{-}\mathrm{str}}(\mathbb{Q},\mathrm{ad}^{0}(E[p^{k}]))=\mathrm{ord}(\epsilon^{\ell}_{1})-\partial\varphi_{\boldsymbol{\epsilon}^{\ell}}(\infty)$
    where $\partial\varphi_{\boldsymbol{\epsilon}^{\ell}}(\infty)$ is the minimal valuation of the whole Stark system $\boldsymbol{\epsilon}^{\ell}$.
  2. (2)
    If $\boldsymbol{\epsilon}^{\ell}$ is primitive, then
    $\mathrm{length}_{\mathbb{Z}_{p}}\mathrm{Sel}_{\ell\textrm{-}\mathrm{str}}(\mathbb{Q},\mathrm{ad}^{0}(E[p^{k}]))=\mathrm{ord}(\epsilon^{\ell}_{1}).$
  3. (3)
    $\boldsymbol{\epsilon}^{\ell}$ is primitive if and only if $\partial\varphi_{\boldsymbol{\epsilon}^{\ell}}(\infty)=0$.
Proof.
See [MR16, Thm. 8.7]. In fact, the module structure of $\mathrm{Sel}_{\ell\textrm{-}\mathrm{str}}(\mathbb{Q},\mathrm{ad}^{0}(E[p^{k}]))$ is also determined by $\boldsymbol{\epsilon}^{\ell}$. ∎
5.3. The key computation
We keep the choices of $\ell$ and $n$ in (5.1) from now on. We explicitly compute
$\Psi_{n,n/q_{s}}\left(c(\ell)\wedge\bigwedge^{s}_{i=1}c(q_{i})\otimes\bigwedge^{s}_{i=1}\psi_{q_{i}}\right)$
where $c(\ell)$ and $c(q_{i})$’s are Flach’s zeta elements in Theorem 2.2 and $\psi_{q_{i}}:\mathrm{H}^{1}_{/f}(\mathbb{Q}_{q_{i}},\mathrm{Sym}^{2}(E[p^{k}]))\simeq\mathbb{Z}/p^{k}\mathbb{Z}$ is the map in Lemma 4.1. Here we use the same notation $c(\ell)$ and $c(q_{i})$ for their image in $\mathrm{H}^{1}(\mathbb{Q},\mathrm{Sym}^{2}(E[p^{k}]))$. Note that $c(\ell)$ and all $c(q_{i})$’s are linearly independent due to their local conditions in Theorem 2.2. We first apply [MR16, Prop. A.1] to our setting. Consider exact sequence
${\mathrm{Sel}_{\mathcal{F}^{\ell n/q_{s}}_{\mathrm{BK}}}(\mathbb{Q},\mathrm{Sym}^{2}(E[p^{k}]))\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathrm{Sel}_{\mathcal{F}^{\ell n}_{\mathrm{BK}}}(\mathbb{Q},\mathrm{Sym}^{2}(E[p^{k}]))\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\psi_{q_{s}}\circ\mathrm{loc}^{s}_{q_{s}}}$${\mathbb{Z}/p^{k}\mathbb{Z}.}$
Then there exists a unique map
$\widehat{\psi_{q_{s}}\circ\mathrm{loc}^{s}_{q_{s}}}:\bigwedge^{s+1}\mathrm{Sel}_{\mathcal{F}^{\ell n}_{\mathrm{BK}}}(\mathbb{Q},\mathrm{Sym}^{2}(E[p^{k}]))\to\mathbb{Z}/p^{k}\mathbb{Z}\otimes\bigwedge^{s}\mathrm{Sel}_{\mathcal{F}^{\ell n/q_{s}}_{\mathrm{BK}}}(\mathbb{Q},\mathrm{Sym}^{2}(E[p^{k}]))$
such that the composition
${\bigwedge^{s+1}\mathrm{Sel}_{\mathcal{F}^{\ell n}_{\mathrm{BK}}}(\mathbb{Q},\mathrm{Sym}^{2}(E[p^{k}]))\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\widehat{\psi_{q_{s}}\circ\mathrm{loc}^{s}_{q_{s}}}}$${\mathbb{Z}/p^{k}\mathbb{Z}\otimes\bigwedge^{s}\mathrm{Sel}_{\mathcal{F}^{\ell n/q_{s}}_{\mathrm{BK}}}(\mathbb{Q},\mathrm{Sym}^{2}(E[p^{k}]))\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathbb{Z}/p^{k}\mathbb{Z}\otimes\bigwedge^{s}\mathrm{Sel}_{\mathcal{F}^{\ell n}_{\mathrm{BK}}}(\mathbb{Q},\mathrm{Sym}^{2}(E[p^{k}]))}$
is given by
$c(\ell)\wedge\bigwedge_{i=1}^{s}c(q_{i})\mapsto\psi_{q_{s}}\circ\mathrm{loc}^{s}_{q_{s}}(c(\ell))\cdot\bigwedge_{i=1}^{s}c(q_{i})+\sum_{i=1}^{s}\left((-1)^{i}\cdot\psi_{q_{s}}\circ\mathrm{loc}^{s}_{q_{s}}(c(q_{i}))\cdot c(\ell)\wedge\bigwedge^{s}_{\begin{subarray}{c}j=1\\ j\neq i\end{subarray}}c(q_{j})\right),$
and the image of $\widehat{\psi_{q_{s}}\circ\mathrm{loc}^{s}_{q_{s}}}$ is the image of $\psi_{q_{s}}\circ\mathrm{loc}^{s}_{q_{s}}\left(\mathrm{Sel}_{\mathcal{F}^{\ell n}_{\mathrm{BK}}}(\mathbb{Q},\mathrm{Sym}^{2}(E[p^{k}]))\right)\otimes\bigwedge^{s}\mathrm{Sel}_{\mathcal{F}^{\ell n/q_{s}}_{\mathrm{BK}}}(\mathbb{Q},\mathrm{Sym}^{2}(E[p^{k}]))$ in $\mathbb{Z}/p^{k}\mathbb{Z}\otimes\bigwedge^{s}\mathrm{Sel}_{\mathcal{F}^{\ell n/q_{s}}_{\mathrm{BK}}}(\mathbb{Q},\mathrm{Sym}^{2}(E[p^{k}]))$.
By the local conditions of Flach’s zeta elements in Theorem 2.2, we have
$\psi_{q_{s}}\circ\mathrm{loc}^{s}_{q_{s}}(c(\ell))\cdot\bigwedge_{i=1}^{s}c(q_{i})+\sum_{i=1}^{s}\left((-1)^{i}\cdot\psi_{q_{s}}\circ\mathrm{loc}^{s}_{q_{s}}(c(q_{i}))\cdot c(\ell)\wedge\bigwedge^{s}_{\begin{subarray}{c}j=1\\ j\neq i\end{subarray}}c(q_{j})\right)$
$=(-1)^{s}\cdot\psi_{q_{s}}\circ\mathrm{loc}^{s}_{q_{s}}(c(q_{s}))\cdot c(\ell)\wedge\bigwedge^{s-1}_{i=1}c(q_{i}).$
Following the concrete description of $\Psi_{n,m}$ in [MR16, pp. 161], we have
(5.2) $\Psi_{n,n/q_{s}}\left(c(\ell)\wedge\bigwedge^{s}_{i=1}c(q_{i})\otimes\bigwedge^{s}_{i=1}\psi_{q_{i}}\right)=(-1)^{s}\cdot\psi_{q_{s}}\circ\mathrm{loc}^{s}_{q_{s}}(c(q_{s}))\cdot c(\ell)\wedge\bigwedge^{s-1}_{i=1}c(q_{i})\otimes\bigwedge^{s-1}_{i=1}\psi_{q_{i}}.$
The computation of the image under $\Psi_{n,n/q_{i}}$ is identical for every $i$ possibly except the sign. By Proposition 5.2, we also have
(5.3) $\Psi_{n,1}=\Psi_{q_{1},1}\circ\Psi_{q_{1}\cdot q_{2},q_{1}}\circ\cdots\circ\Psi_{n/q_{s},n/(q_{s-1}\cdot q_{s})}\circ\Psi_{n,n/q_{s}}.$
5.4. A Stark system of Gauss sum type and its extension
We keep the choices of $\ell$ and $n$ in (5.1) as before. Define
$\epsilon^{\mathrm{Flach}?,\ell}_{n}:=(-1)^{\frac{s(s+1)}{2}}\cdot c(\ell)\wedge\bigwedge^{s}_{i=1}c(q_{i})\otimes\bigwedge^{s}_{i=1}\psi_{q_{i}}\in Y_{n}$
so that
$\Psi_{n,1}(\epsilon^{\mathrm{Flach}?,\ell}_{n})=\left(\prod_{i=1}^{s}\psi_{q_{i}}\circ\mathrm{loc}^{s}_{q_{i}}(c(q_{i}))\right)\cdot c(\ell).$
This formula follows from the iteration of (5.2) via (5.3).
For $m$ dividing $n$, we define
$\epsilon^{\mathrm{Flach}?,\ell}_{m}:=\Psi_{n,m}(\epsilon^{\mathrm{Flach}?,\ell}_{n}).$
By Definition 5.3, $\left\{\epsilon^{\mathrm{Flach}?,\ell}_{m}\in Y_{m}:m|n\right\}$ forms a finite Stark system of core rank one, i.e. a Stark system for Selmer data $(\mathrm{Sym}^{2}(E[p^{k}]),\mathcal{F}^{\ell}_{\mathrm{BK}},\mathcal{P}^{n})$ where $\mathcal{P}^{n}$ is the (finite!) set of the primes dividing $n$.
Let $m$ is a divisor of $n$ and $G_{m}=\otimes_{q|m}\mathrm{Gal}(\mathbb{Q}(\zeta^{(p)}_{q})/\mathbb{Q})$ where $\mathbb{Q}(\zeta^{(p)}_{q})$ is the maximal $p$-subextension of $\mathbb{Q}$ in $\mathbb{Q}(\zeta_{q})$. Recall the map in [MR16, §$\S$12]
$\Pi_{m}:Y_{m}\to\mathrm{Sel}_{\mathcal{F}^{\ell}_{\mathrm{BK}}(m)}(\mathbb{Q},\mathrm{Sym}^{2}(E[p^{k}]))\otimes G_{m}$
where $\mathcal{F}^{\ell}_{\mathrm{BK}}(m)$ is the $\ell$-relaxed and $m$-transverse Bloch–Kato Selmer structure. Then [MR16, Prop. 12.3] implies that
$\left\{(-1)^{\nu(m)}\cdot\Pi_{m}(\epsilon^{\mathrm{Flach}?,\ell}_{m}):m|n\right\}$
forms a finite Kolyvagin system for $\left(\mathrm{Sym}^{2}(E[p^{k}]),\mathcal{F}^{\ell}_{\mathrm{BK}},\mathcal{P}^{n}\right)$. Thanks to the core rank one property recalled in §$\S$3.7, the finite Kolyvagin system extends to the rank one Kolyvagin system $\boldsymbol{\kappa}^{\mathrm{Flach}?,\ell}$ for $\left(\mathrm{Sym}^{2}(E[p^{k}]),\mathcal{F}^{\ell}_{\mathrm{BK}},\mathcal{P}^{\mathrm{Flach},(\ell)}_{k}\right)$ as in the case of Gauss sum Kolyvagin systems [MR04, Rem. 6.3.3].
Remark 5.6.
If $n\in\mathcal{P}^{\mathrm{Flach},(\ell)}_{k}$ with $\mathrm{Sel}_{\mathcal{F}_{\mathrm{BK},\ell n}}(\mathbb{Q},\mathrm{ad}^{0}(E[p^{k}]))=0$, then $\kappa^{\mathrm{Flach}?,\ell}_{n}$ is uniquely determined by $\kappa^{\mathrm{Flach}?,\ell}_{m}$ for $m$ properly dividing $n$ [MR04, Rem. 3.1.9] where $\mathcal{F}_{\mathrm{BK},\ell n}$ is the $\ell n$-strict Bloch–Kato Selmer structure. From this point of view, the extension of a finite Kolyvagin system is not a miracle and it shows the strength of the rigidity of Kolyvagin systems.
The modules of Stark and Kolyvagin systems over $\mathbb{Z}/p^{k}\mathbb{Z}$ are isomorphic as free $\mathbb{Z}/p^{k}\mathbb{Z}$-modules of rank one (Theorem 5.4 and [MR16, Rem. 11.5 and Thm. 12.4]). By using this isomorphism, the finite Stark system also extends to the rank one Stark system $\boldsymbol{\epsilon}^{\mathrm{Flach}?,\ell}$ for $\left(\mathrm{Sym}^{2}(E[p^{k}]),\mathcal{F}^{\ell}_{\mathrm{BK}},\mathcal{P}^{\mathrm{Flach},(\ell)}_{k}\right)$.
5.5. The primitive normalization
We construct the primitive Stark system $\boldsymbol{\epsilon}^{\mathrm{Flach},\ell}$ from $\boldsymbol{\epsilon}^{\mathrm{Flach}?,\ell}$ and compute the constant multiple in Theorem 1.1.
By global Poitou–Tate duality [Rub00, Thm. 1.7.3], we have exact sequence
(5.4) ${\mathrm{Sel}_{\mathcal{F}_{\mathrm{BK},\ell}}(\mathbb{Q},\mathrm{ad}^{0}(E[p^{k}]))\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathrm{Sel}_{\mathcal{F}_{\mathrm{BK}}}(\mathbb{Q},\mathrm{ad}^{0}(E[p^{k}]))\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\left(\dfrac{\mathrm{H}^{1}_{/f}(\mathbb{Q}_{\ell},\mathrm{Sym}^{2}(E[p^{k}]))}{\mathrm{loc}^{s}_{\ell}\left(\mathrm{Sel}_{\mathcal{F}^{\ell}_{\mathrm{BK}}}(\mathbb{Q},\mathrm{Sym}^{2}(E[p^{k}]))\right)}\right)^{\vee}}$
where $(-)^{\vee}$ is the Pontryagin dual. We also have
$\kappa^{\mathrm{Flach}?,\ell}_{1}$ $=\epsilon^{\mathrm{Flach}?,\ell}_{1}$
$=\left(\prod_{i=1}^{s}\psi_{q_{i}}\circ\mathrm{loc}^{s}_{q_{i}}(c(q_{i}))\right)\cdot c(\ell)$
$\in\mathrm{Sel}_{\mathcal{F}^{\ell}_{\mathrm{BK}}}(\mathbb{Q},\mathrm{Sym}^{2}(E[p^{k}]))$
$\simeq\mathbb{Z}/p^{k}\mathbb{Z}\oplus\mathrm{Sel}_{\mathcal{F}_{\mathrm{BK},\ell}}(\mathbb{Q},\mathrm{ad}^{0}(E[p^{k}]))$
where the last isomorphism is non-canonical and follows from [MR16, Cor. 3.5.(i)]. By [MR04, Thm. 4.4.1], we know $\langle\kappa^{\mathrm{Flach}?,\ell}_{1}\rangle\subseteq\mathbb{Z}/p^{k}\mathbb{Z}$ in the above decomposition. Thus, we have equality
(5.5) $\mathrm{ord}_{p}(c(\ell))+\mathrm{length}_{\mathbb{Z}_{p}}\left(\dfrac{\mathrm{H}^{1}_{/f}(\mathbb{Q}_{\ell},\mathrm{Sym}^{2}(E[p^{k}]))}{\mathrm{loc}^{s}_{p}\left(\mathrm{Sel}_{\mathcal{F}^{\ell}_{\mathrm{BK}}}(\mathbb{Q},\mathrm{Sym}^{2}(E[p^{k}]))\right)}\right)^{\vee}=\mathrm{ord}_{p}(\mathrm{loc}^{s}_{p}(c(\ell)))$
where $\mathrm{ord}_{p}(a)$ means the $p$-divisibility index of $a$ in the natural module containing $a$.
Write $\lambda=\mathrm{length}_{\mathbb{Z}_{p}}\mathrm{Sel}_{\mathcal{F}_{\mathrm{BK}}}(\mathbb{Q},\mathrm{ad}^{0}(E[p^{k}]))$ for convenience. Consider the tautological bound
$\mathrm{length}_{\mathbb{Z}_{p}}\mathrm{Sel}_{\mathcal{F}_{\mathrm{BK}}}(\mathbb{Q},\mathrm{ad}^{0}(E[p^{k}]))=\lambda=\mathrm{ord}_{p}\left(\dfrac{p^{\lambda}}{\psi_{\ell}\circ\mathrm{loc}^{s}_{\ell}(c(\ell))}\cdot\mathrm{loc}^{s}_{\ell}(c(\ell))\right)$
where $\psi_{\ell}:\mathrm{H}^{1}_{/f}(\mathbb{Q}_{\ell},\mathrm{Sym}^{2}(E[p^{k}]))\simeq\mathbb{Z}/p^{k}\mathbb{Z}$ is the map in Lemma 4.1. By (5.4) and (5.5), we also obtain the exact bound the $\ell$-strict one
(5.6) $\mathrm{length}_{\mathbb{Z}_{p}}\mathrm{Sel}_{\mathcal{F}_{\mathrm{BK},\ell}}(\mathbb{Q},\mathrm{ad}^{0}(E[p^{k}]))=\mathrm{ord}_{p}\left(\dfrac{p^{\lambda}}{\psi_{\ell}\circ\mathrm{loc}^{s}_{\ell}(c(\ell))}\cdot c(\ell)\right).$
We apply Theorem 5.5 to $\boldsymbol{\epsilon}^{\mathrm{Flach}?,\ell}$ and compare it with (5.6). Then we obtain
(5.7) $\begin{split}&\mathrm{length}_{\mathbb{Z}_{p}}\mathrm{Sel}_{\mathcal{F}_{\mathrm{BK},\ell}}(\mathbb{Q},\mathrm{ad}^{0}(E[p^{k}]))\\ &=\mathrm{ord}_{p}\left(\left(\prod_{i=1}^{s}\psi_{q_{i}}\circ\mathrm{loc}^{s}_{q_{i}}(c(q_{i}))\right)\cdot c(\ell)\right)+\mathrm{ord}_{p}\left(\dfrac{p^{\lambda}}{\psi_{\ell}\circ\mathrm{loc}^{s}_{\ell}(c(\ell))}\right)-\mathrm{ord}_{p}\left(\prod_{i=1}^{s}\psi_{q_{i}}\circ\mathrm{loc}^{s}_{q_{i}}(c(q_{i}))\right),\end{split}$
so the minimal valuation of the whole Stark system $\boldsymbol{\epsilon}^{\mathrm{Flach}?,\ell}$ is
$\mathrm{ord}_{p}\left(\dfrac{\psi_{\ell}\circ\mathrm{loc}^{s}_{\ell}(c(\ell))}{p^{\lambda}}\cdot\prod_{i=1}^{s}\psi_{q_{i}}\circ\mathrm{loc}^{s}_{q_{i}}(c(q_{i}))\right).$
Thus, we define the primitive Flach–Stark system by
$\boldsymbol{\epsilon}^{\mathrm{Flach},\ell}=\dfrac{p^{\lambda}}{\psi_{\ell}\circ\mathrm{loc}^{s}_{\ell}(c(\ell))}\cdot\left(\prod_{i=1}^{s}\dfrac{1}{\psi_{q_{i}}\circ\mathrm{loc}^{s}_{q_{i}}(c(q_{i}))}\right)\cdot\boldsymbol{\epsilon}^{\mathrm{Flach}?,\ell}$
in the module of Stark systems, and we have explicit formulas
$\epsilon^{\mathrm{Flach},\ell}_{1}$ $=\dfrac{p^{\lambda}}{\psi_{\ell}\circ\mathrm{loc}^{s}_{\ell}(c(\ell))}\cdot c(\ell),$
$\epsilon^{\mathrm{Flach},\ell}_{n^{\prime}}$ $=\dfrac{(-1)^{\frac{s^{\prime}(s^{\prime}+1)}{2}}\cdot p^{\lambda}}{\psi_{\ell}\circ\mathrm{loc}^{s}_{\ell}(c(\ell))}\cdot\left(\prod_{i=1}^{s^{\prime}}\dfrac{1}{\psi_{q^{\prime}_{i}}\circ\mathrm{loc}^{s}_{q^{\prime}_{i}}(c(q^{\prime}_{i}))}\right)\cdot c(\ell)\wedge\bigwedge^{s^{\prime}}_{i=1}c(q^{\prime}_{i})\otimes\bigwedge^{s^{\prime}}_{i=1}\psi_{q^{\prime}_{i}}$
for every $n^{\prime}=q^{\prime}_{1}\cdot q^{\prime}_{2}\cdot\cdots\cdot q^{\prime}_{s^{\prime}}\in\mathcal{N}^{\mathrm{Flach},(\ell)}_{k}$ with $\nu(n^{\prime})=s^{\prime}\geq 1$. Note that our choice of $n\in\mathcal{N}^{\mathrm{Flach},(\ell)}_{k}$ in (5.1) was arbitrary (Remark 5.1).
In particular, the constant multiple in Theorem 1.1 is $\dfrac{p^{\lambda}}{\psi_{\ell}\circ\mathrm{loc}^{s}_{\ell}(c(\ell))}=\dfrac{p^{\lambda}}{\mathrm{deg}(\phi)}$ by construction and by the explicit reciprocity law (2.1). Therefore, the Bloch–Kato conjecture (Corollary 1.4.(2)) is equivalent to showing that this constant multiple is a $p$-adic unit.
Remark 5.7.
Although each $\dfrac{c(q^{\prime}_{i})}{\psi_{q^{\prime}_{i}}\circ\mathrm{loc}^{s}_{q^{\prime}_{i}}(c(q^{\prime}_{i}))}$ makes sense if and only if $\mathrm{loc}^{s}_{q^{\prime}_{i}}$ is surjective, $\epsilon^{\mathrm{Flach},\ell}_{n^{\prime}}$ makes sense without assuming the surjectivity of each $\mathrm{loc}^{s}_{q^{\prime}_{i}}$. In fact, the possible failure of the surjectivity of these localization maps makes the Stark system argument non-vacuous.
6. Proofs of corollaries
6.1. Proof of Corollary 1.2
This follows from the tautological bound in §5.5 and the constant multiple $\dfrac{p^{\lambda}}{\psi_{\ell}\circ\mathrm{loc}^{s}_{\ell}(c(\ell))}=\dfrac{p^{\lambda}}{\mathrm{deg}(\phi)}$.
6.2. Proof of Corollary 1.4
The equivalence between (1) and (2) follows from the last argument in §$\S$5.5. The equivalence between the Bloch–Kato conjecture (2) and the modularity lifting theorem (3) follows from Wiles’ numerical criterion [DDT97, Thm. 5.3]. See also [DFG04].
6.3. Proof of Corollary 1.5
The assumption says that
$\mathrm{ord}_{p}\left(\dfrac{p^{\lambda}}{\psi_{\ell}\circ\mathrm{loc}^{s}_{\ell}(c(\ell))}\cdot c(\ell)\right)=\mathrm{ord}_{p}\left(c(\ell)\right)=0.$
Thus, (5.6) implies $\mathrm{length}_{\mathbb{Z}_{p}}\mathrm{Sel}_{\mathcal{F}_{\mathrm{BK},\ell}}(\mathbb{Q},\mathrm{ad}^{0}(E[p^{k}]))=0$. By global duality (5.4), we have
$\mathrm{Sel}_{\mathcal{F}_{\mathrm{BK}}}(\mathbb{Q},\mathrm{ad}^{0}(E[p^{k}]))\simeq\left(\dfrac{\mathrm{H}^{1}_{/f}(\mathbb{Q}_{\ell},\mathrm{Sym}^{2}(E[p^{k}]))}{\mathrm{loc}^{s}_{\ell}\left(\mathrm{Sel}_{\mathcal{F}^{\ell}_{\mathrm{BK}}}(\mathbb{Q},\mathrm{Sym}^{2}(E[p^{k}]))\right)}\right)^{\vee}$
and $\mathrm{H}^{1}_{/f}(\mathbb{Q}_{\ell},\mathrm{Sym}^{2}(E[p^{k}]))\simeq\mathbb{Z}/p^{k}\mathbb{Z}$ by Lemma 4.1, so the proof is complete.
Acknowledgement
We thank Francesc Castella, Henri Darmon, Matthias Flach, Minhyong Kim, Masato Kurihara, Tom Weston, and Christopher Skinner for their interests in our work. We are benefited from the discussion with Ashay Burungale, Haruzo Hida, Antonio Lei, Gyujin Oh, and Ryotaro Sakamoto. We would like to thank Marco Sangiovanni Vincentelli for kindly sharing [SS] with us and for his interest on the cyclicity result.